OpenStudy (anonymous):

The function f(x) = x². The graph of g(x) is f(x) translated to the right 3 units and down 3 units. What is the function rule for g(x)?

OpenStudy (anonymous):

@precal

OpenStudy (precal):

right is -3 and down is -3

OpenStudy (anonymous):

g(x) = -3x+2

OpenStudy (anonymous):

g(x) = f(x−3) − 3 g(x) = (x−3)² − 3

OpenStudy (precal):

what happened to the second power

OpenStudy (precal):

I don't want your choices I want you to tell me what you think the answer is

OpenStudy (anonymous):

i did, this is my awnsr g(x) = f(x−3) − 3 g(x) = (x−3)² − 3

OpenStudy (precal):

you were given a function so I would use the function

OpenStudy (precal):

did you figure out which one it is?

OpenStudy (anonymous):

no-.-

OpenStudy (precal):